OpenStudy (anonymous):

Which of the following words does not mean “subtract"? A. difference B. quotient C. less than D. decreased by

OpenStudy (anonymous):

B

OpenStudy (anonymous):

b duhh

OpenStudy (anonymous):

Write an algebraic equation for the following. The length, l, of a rectangle is sixteen less than seven times the width, w. A. l = 16 – 7w B. l – 16 = 7w C. l = 7w – 16 D. 16 – l = 7w

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b aww u beat me

OpenStudy (anonymous):

Two redwood trees are the same height. A third tree's height is x feet. The first tree is 50 feet more than three times the height of the third tree. The second tree is 100 feet taller than twice the height of the third tree. What is the height, in feet, of the third tree? A. 150 B. 50 C. 200 D. 100

OpenStudy (anonymous):

a

OpenStudy (anonymous):

The number of pizzas eaten at the class party is equal to one-fourth the number of students, plus two more pizzas for the adults. If nine pizzas are eaten, how many students are there? A. 28 B. 2 C. 44 D. 36

OpenStudy (anonymous):

1/4a + 2 = 9

OpenStudy (anonymous):

a = 28

OpenStudy (anonymous):

Five more than a number tripled is half the number. What is the number? A. 6 B. –2 C. –1 D. 1

OpenStudy (anonymous):

a

OpenStudy (anonymous):

omg every time lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

5 + 3a = 1/2a

OpenStudy (anonymous):

a = -2